OpenStudy (anonymous):

What is the sine value of 5 pi over 3?

OpenStudy (anonymous):

negative square root 3 over 2 square root 3 over 2 negative 1 over 2 1 over 2

OpenStudy (anonymous):

5pi/3 = (5*180)/3 degrees = 300 degrees sin(300 degrees) = \[\frac{ -\sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

wait it's negative?

OpenStudy (anonymous):

The fraction is, yes

OpenStudy (anonymous):

okay thanks! can you help me with one other thing?

OpenStudy (anonymous):

sure